# Complete asymptotics in the formation of quiescent big bang singularities

### v1

@@ 1. Introduction
Studying the stability of big bang formation is an alternative approach which has been very active in the last 5–10 years. The results began with the work of Rodnianski and Speck, demonstrating stability of spatially homogeneous, isotropic and spatially flat solutions in the Einstein-scalar field and the Einstein-stiff fluid setting; see [47, 48]. Speck demonstrates a similar result in the case of $\mathbb{S}^{3}$-spatial topology; see [50]. In [49], the authors consider the vacuum equations in higher dimensions in situations with intermediate anisotropy; the conditions are more general than requiring proximity to isotropy, but they are still quite far from covering the full subcritical range in which stability is expected. In [18], Fajman and Urban prove a stability result in the case of spatial local homogeneity and isotropy and negative spatial curvature. This result was later generalised to allow the presence of Vlasov matter in [19]. Finally, Fournodavlos, Rodnianski and Speck demonstrate stable big bang formation for the full subcritical range of Kasner-scalar field spacetimes in [21]. This is a very important result for many reasons. First, because it covers the maximal range in which one can obtain stability. Second, because the authors introduce a new formulation of the equations using a Fermi-Walker propagated frame which has turned out to be very fruitful.
@@ 1.1. The Einstein–nonlinear scalar field equations
Due to arguments going back to the work of Yvonne Choquet-Bruhat and Robert Geroch [23], there is a unique maximal globally hyperbolic development associated with initial data; see also [40, Chapter 23] for a textbook presentation.
@@
Let $(M,g,\varphi)$ be a solution to (1) and let $\Sigma$ be a spacelike hypersurface in $(M,g)$ with future directed unit normal $U$. Define $h$ and $k$ to be the associated induced first and second fundamental forms. Let, moreover, $\phi:=\varphi|_{\Sigma}$ and $\psi:=U(\varphi)|_{\Sigma}$. Then $(\Sigma,h,k,\phi,\psi)$ are initial data for the Einstein–nonlinear scalar field equations; see, e.g., [38, Chapter 13].
@@ 1.2. Initial data on the singularity
It is possible to omit the nondegeneracy requirement; i.e., the requirement that the eigenvalues be distinct. However, we are mainly interested in the nondegenerate setting in this article. It is also possible to omit the requirement that (4) hold. However, it is then necessary to introduce separate conditions on the structure coefficients associated with an eigenframe of $\mathring{\mathcal{K}}$; see, e.g., [43, Definition 10, p. 9]. Moreover, the solutions inducing such initial data are expected to be unstable. Since we are here interested in the asymptotics induced by solutions arising in stability results, such a generalisation is therefore not of great interest here.
@@
In $3+1$-dimensions, the step from data on the singularity to corresponding developments is discussed in [22]. In particular, due to [22, Theorem 1.21, p. 10], initial data on the singularity give rise to unique (up to isometry) maximal globally hyperbolic developments with a local Gaussian foliation close to the singularity. Next, we turn to the opposite direction. In particular, we prove that the general condition on initial data introduced in [32], ensuring that the corresponding maximal globally hyperbolic developments have crushing CMC singularities with curvature blow up, in fact yields solutions inducing data on the singularity along the CMC foliation.
@@ 1.3. The main result
In comparison with [32, Definition 1, p. 5], we have here omitted the requirement that $V$ be non-negative. We include this requirement explicitly in Theorem 11 instead.
@@
The constant $\kappa$ is given as follows. In the main theorem of [32], i.e. [32, Theorem 12, pp. 6–7], the constants $k_{0}$ and $k_{1}$ are introduced. Choosing $k_{0}$ by demanding that equality hold in [32, (12a), p. 6]; and choosing $k_{1}$ to be the smallest integer such that [32, (12b), p. 6] holds for this choice of $k_{0}$, the integer $\kappa$ is given by $\kappa=k_{1}+2$.
@@
The requirement that $V$ be non–negative can be relaxed; cf. [32, Remark 16, p. 7].
@@ 1.4. The FRS equations
One topic of interest in this article is solutions to the FRS equations, introduced in [21]. The purpose of the present section is to recall the slight reformulation of these equations introduced in [32, Theorem 12, pp. 5–6]. These equations arise from the Einstein–nonlinear scalar field equations by assuming a CMC foliation, a vanishing shift vector field, and using a Fermi-Walker transported frame. The underlying assumptions are the following. Let $(\Sigma,h_{\mathrm{ref}})$ be a closed parallelisable Riemannian manifold of dimension $n$; let $\{E_{i}\}_{i=1}^{n}$ be a smooth global orthonormal frame with respect to $h_{\mathrm{ref}}$ with dual frame $\{\eta^{i}\}_{i=1}^{n}$; and let $V\in C^{\infty}(\mathbb{R})$. The manifold on which we wish to solve the equations is of the form $M:=\Sigma\times\mathcal{I}$ for some open interval $\mathcal{I}\subset(0,\infty)$ and the basic variables are $N$, $e_{I}^{i}$, $\omega^{I}_{i}$, $\gamma_{IJK}$, $k_{IJ}$ and $\varphi$. Moreover, we assume $N>0$ on $M$. As clarified in the statement of [32, Proposition 68, pp. 23–24], the geometric interpretation of these variables is the following. Let $e_{0}:=N^{-1}\partial_{t}$, where $\partial_{t}$ signifies differentiation with respect to the second variable in $M\times\mathcal{I}$; $e_{I}:=e_{I}^{i}E_{i}$; and $\omega^{I}:=\omega^{I}_{i}\eta^{i}$. Then $\{e_{\alpha}\}_{\alpha=0}^{n}$ is a time oriented orthonormal frame for a Lorentz metric $g$ on $M$ and if $h$ denotes the induced metric on the constant-$t$ hypersurfaces $\Sigma_{t}:=\Sigma\times\{t\}$, then $\{e_{I}\}_{I=1}^{n}$ is an orthonormal frame with respect to $h$ with dual frame $\{\omega^{I}\}_{I=1}^{n}$. Moreover,
@@
are the induced second fundamental form on $\Sigma_{t}$ and the structure coefficients of the frame $\{e_{I}\}_{I=1}^{n}$, respectively, and the mean curvature satisfies $k_{II}=t^{-1}$ (here and below, we sum over repeated capital Latin indices, even if both are subscripts). Given the above underlying assumptions, the equations, see [32, Proposition 68, pp. 23–24], are the following, using the conventions introduced in Subsection A.2:
@@
It is convenient to recall that combining (17) and (19) yields [32, (85), p. 32]:
@@
In the above equations and in what follows, it is useful to recall [32, Lemma 70, p. 25]:
@@ 1.5. Higher order derivative estimates for solutions to the FRS equations
Even though our main application in the present paper is to derive complete asymptotics for the solutions constructed in [32], it is of interest to derive energy estimates using minimal assumptions. In this way, the conclusions might be useful in other contexts. The main assumptions are the following.
@@ 1.6. Deriving asymptotics, given higher order derivative bounds
To begin with, (30a) is, e.g., an immediate consequence of [42, (269), p. 65], keeping in mind that
@@
$\mathcal{P}$ is defined by [42, (17), p. 8] and the fact that $T$ is given by (2). Next, (30b) follows from a straightforward computation and (31) follow from [38, Chapter 13].
Turning to the other direction, (1b) is an immediate consequence of (30b). Due to (31) the normal–normal and the normal–tangential components of Einstein’s equations hold. Next, combining [42, (257), p. 63] and [42, (269), p. 65] with (30a), it is clear that the spatial–spatial components of the spacetime Ricci tensor equal what they would if Einstein’s equations would hold. Combining this information with the fact that the normal–normal and the normal–tangential components of Einstein’s equations hold, it can be deduced that (1a) holds. ∎
@@
(36c) $\|N-1\|_{C^{k_{0}}(\Sigma)}+t^{1-\delta}\|\vec{e}(\ln N)\|_{C^{k_{0}}(\Sigma)}+t^{1-\delta}\|\vec{e}\varphi\|_{C^{k_{0}}(\Sigma)}$ $\leq\mathbf{c}_{0},$
(36d) $\|t\Theta-1\|_{C^{0}(\Sigma)}+\|N-1\|_{C^{0}(\Sigma)}$ $\leq\tfrac{1}{2}$
@@ 1.7. Outline of the arguments
The main result of this article is Theorem 11. However, Theorem 11 is an immediate consequence of Theorems 18 and 22. Moreover, the latter theorems have an independent interest in their own right. In particular, it is to be expected that Theorem 22 can be used to obtain data on the singularity in other settings. For these reasons, we here focus on a description of the arguments required to prove Theorems 18 and 22.
@@
Following the results of [32], it makes sense to set as our starting point some $C^{k_{0}}$ bounds on a Fermi-Walker frame and on the components of geometric quantities in terms of this frame, along with $C^{k_{0}}$ convergence of the eigenvalues of $\mathcal{K}$, of $\Phi$ and of $\Psi$. These are precisely the estimates (36) and (35). Note that these estimates correspond to the conclusions obtained in [32] with two exceptions. First, since we do not assume the CMC condition, we need a way to control the mean curvature. This is done through the quantity $\Theta=N\theta$, which is assumed to be asymptotic to $t^{-1}$ as $t\downarrow 0$. Second, in [32], the authors obtain convergence of the lapse, $N\to 1$ in $C^{k_{0}}$ as $t\downarrow 0$. By contrast, here we only assume that N1N-1 is bounded in $C^{k_{0}}$, and not too large in $C^{0}$. However, we point out that if the CMC condition is assumed, then the assumption that $t\Theta-1\to 0$ becomes $N-1\to 0$. We begin by deriving some basic consequences of the assumptions in Subsection 3.1.
@@
Proof of the main result. Once we have Theorems 18 and 22 at our disposal, we can combine them with the arguments in [32] in order to deduce that Theorem 11 holds. The details are to be found in Section 4.
@@ 1.8. Acknowledgements
This research was funded by the Swedish Research Council (Vetenskapsrådet), dnr. 2022-03053.
@@ 2.5. Estimating the lapse function
As in the proof of [32, Lemma 110, p. 52], applying $E_{\mathbf{I}}$ to (20) and then multiplying with $E_{\mathbf{I}}(N-1)$ yields
@@
(61) $\begin{split}&-\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}e_{I}(N-1)+t^{-2}\big(E_{\mathbf{I}}(N-1)\big)^{2}\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}e_{I}(N-1)[e_{I},E_{\mathbf{I}}](N-1)-[E_{\mathbf{I}},e_{I}]e_{I}(N-1)E_{\mathbf{I}}(N-1)\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}(N-1)\mathrm{div}_{h_{\mathrm{ref}}}e_{I}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ ={}&\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}\big(\gamma_{JII}e_{J}(N)-N\big\{t^{-2}-k_{IJ}k_{IJ}-e_{0}(\varphi)e_{0}(\varphi)+\tfrac{2}{n-1}V\circ\varphi\big\}\big)E_{\mathbf{I}}(N-1)\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
Next, note that, due to (39),

### v2

# arXiv:2602.02373v2[gr-qc] 02 Mar 2026

@@ 1. Introduction
Studying the stability of big bang formation is an alternative approach which has been very active in the last 5–10 years. The results began with the work of Rodnianski and Speck, demonstrating stability of spatially homogeneous, isotropic and spatially flat solutions in the Einstein-scalar field and the Einstein-stiff fluid setting; see [49, 50]. Speck demonstrates a similar result in the case of $\mathbb{S}^{3}$-spatial topology; see [52]. In [51], the authors consider the vacuum equations in higher dimensions in situations with intermediate anisotropy; the conditions are more general than requiring proximity to isotropy, but they are still quite far from covering the full subcritical range in which stability is expected. In [19], Fajman and Urban prove a stability result in the case of spatial local homogeneity and isotropy and negative spatial curvature. This result was later generalised to allow the presence of Vlasov matter and any sign of the sectional curvature in [20]. Finally, Fournodavlos, Rodnianski and Speck demonstrate stable big bang formation for the full subcritical range of Kasner-scalar field spacetimes in [22]. This is a very important result for many reasons. First, because it covers the maximal range in which one can obtain stability. Second, because the authors introduce a new formulation of the equations using a Fermi-Walker propagated frame which has turned out to be very fruitful. The extension of this result to stability of the big bang singularity for the Einstein-Maxwell-scalar field-Vlasov system in the full strong subcritical range is discussed in [3].
@@ 1.1. The Einstein–nonlinear scalar field equations
Due to arguments going back to the work of Yvonne Choquet-Bruhat and Robert Geroch [24], there is a unique maximal globally hyperbolic development associated with initial data; see also [41, Chapter 23] for a textbook presentation.
@@
Let $(M,g,\varphi)$ be a solution to (1) and let $\Sigma$ be a spacelike hypersurface in $(M,g)$ with future directed unit normal $U$. Define $h$ and $k$ to be the associated induced first and second fundamental forms. Let, moreover, $\phi:=\varphi|_{\Sigma}$ and $\psi:=U(\varphi)|_{\Sigma}$. Then $(\Sigma,h,k,\phi,\psi)$ are initial data for the Einstein–nonlinear scalar field equations; see, e.g., [39, Chapter 13].
@@ 1.2. Initial data on the singularity
It is possible to omit the nondegeneracy requirement; i.e., the requirement that the eigenvalues be distinct. However, we are mainly interested in the nondegenerate setting in this article. It is also possible to omit the requirement that (4) hold. However, it is then necessary to introduce separate conditions on the structure coefficients associated with an eigenframe of $\mathring{\mathcal{K}}$; see, e.g., [45, Definition 10, p. 9]. Moreover, the solutions inducing initial data violating (4) are expected to be unstable. Since we are here interested in the asymptotics induced by solutions arising in stability results, such a generalisation is therefore not of great interest here.
@@
In $3+1$-dimensions, the step from data on the singularity to corresponding developments is discussed in [23]. In particular, due to [23, Theorem 1.21, p. 10], initial data on the singularity give rise to unique (up to isometry) maximal globally hyperbolic developments with a local Gaussian foliation close to the singularity. Next, we turn to the opposite direction. In particular, we prove that the general condition on initial data introduced in [33], ensuring that the corresponding maximal globally hyperbolic developments have crushing CMC singularities with curvature blow up, in fact yields solutions inducing data on the singularity along the CMC foliation.
@@ 1.3. The main result
In comparison with [33, Definition 1, p. 5], we have here omitted the requirement that $V$ be non-negative. We include this requirement explicitly in Theorem 11 instead.
@@
The constant $\kappa$ is given as follows. In the main theorem of [33], i.e. [33, Theorem 12, pp. 6–7], the constants $k_{0}$ and $k_{1}$ are introduced. Choosing $k_{0}$ by demanding that equality hold in [33, (12a), p. 6]; and choosing $k_{1}$ to be the smallest integer such that [33, (12b), p. 6] holds for this choice of $k_{0}$, the integer $\kappa$ is given by $\kappa=k_{1}+2$.
@@
The requirement that $V$ be non–negative can be relaxed; cf. [33, Remark 16, p. 7].
Remark 15.
In [33], the stability of quiescent model solutions, a notion introduced in [33, Definition 30, p. 11], is also discussed; cf. [33, Theorem 33, p. 12] and [33, Subsection 1.6, pp. 13–19]. The background solutions to the Einstein-scalar field equations considered in, e.g., [52, 22, 19, 20] are examples of quiescent model solutions, but the notion is more general than this. There are quiescent model solutions inducing degenerate data on the singularity; i.e., data such that the $\mathring{p}_{I}$ are not all distinct. For this reason, the solutions obtained by appealing to the stability results are not necessarily such that Theorem 11 applies. However, even in the degenerate setting, one does obtain conclusions. Consider a solution obtained by appealing to [33, Theorem 33, p. 12]. Then the first part of the proof of Theorem 11, which simply consists of an application of Theorem 19, still applies. Next, assume that at some point, say $x\in\Sigma$, the eigenvalues $\mathring{p}_{I}(x)$ are all distinct. Then there is a closed ball, say $\bar{B}_{R}(x)$, of radius $R>0$ with respect to the reference metric $h_{\mathrm{ref}}$ such that, for every $y\in\bar{B}_{R}(x)$, the eigenvalues $\mathring{p}_{I}(y)$ are all distinct. Choosing $T>0$, $\delta>0$ and $r>0$ small enough, it can then be ensured that $g$ and $\varphi$, restricted to $M_{\mathrm{loc}}:=\bar{B}_{R}(x)\times(0,T)$, constitute a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions for any $k_{0}$. This means that we can apply Theorem 24 in order to obtain smooth initial data on the singularity in the ball $\bar{B}_{R}(x)$. Moreover, (9) holds with $\Sigma$ replaced by $\bar{B}_{R}(x)$.
@@ 1.4. The FRS equations
One topic of interest in this article is solutions to the FRS equations, introduced in [22]. The purpose of the present section is to recall the slight reformulation of these equations introduced in [33, Theorem 12, pp. 5–6]. These equations arise from the Einstein–nonlinear scalar field equations by assuming a CMC foliation, a vanishing shift vector field, and using a Fermi-Walker transported frame. The underlying assumptions are the following. Let $(\Sigma,h_{\mathrm{ref}})$ be a closed parallelisable Riemannian manifold of dimension $n$; let $\{E_{i}\}_{i=1}^{n}$ be a smooth global orthonormal frame with respect to $h_{\mathrm{ref}}$ with dual frame $\{\eta^{i}\}_{i=1}^{n}$; and let $V\in C^{\infty}(\mathbb{R})$. The manifold on which we wish to solve the equations is of the form $M:=\Sigma\times\mathcal{I}$ for some open interval $\mathcal{I}\subset(0,\infty)$ and the basic variables are $N$, $e_{I}^{i}$, $\omega^{I}_{i}$, $\gamma_{IJK}$, $k_{IJ}$ and $\varphi$. Moreover, we assume $N>0$ on $M$. As clarified in the statement of [33, Proposition 70, pp. 23–24], the geometric interpretation of these variables is the following. Let $e_{0}:=N^{-1}\partial_{t}$, where $\partial_{t}$ signifies differentiation with respect to the second variable in $M\times\mathcal{I}$; $e_{I}:=e_{I}^{i}E_{i}$; and $\omega^{I}:=\omega^{I}_{i}\eta^{i}$. Then $\{e_{\alpha}\}_{\alpha=0}^{n}$ is a time oriented orthonormal frame for a Lorentz metric $g$ on $M$ and if $h$ denotes the induced metric on the constant-$t$ hypersurfaces $\Sigma_{t}:=\Sigma\times\{t\}$, then $\{e_{I}\}_{I=1}^{n}$ is an orthonormal frame with respect to $h$ with dual frame $\{\omega^{I}\}_{I=1}^{n}$. Moreover,
@@
are the induced second fundamental form on $\Sigma_{t}$ and the structure coefficients of the frame $\{e_{I}\}_{I=1}^{n}$, respectively, and the mean curvature satisfies $k_{II}=t^{-1}$ (here and below, we sum over repeated capital Latin indices, even if both are subscripts). Given the above underlying assumptions, the equations, see [33, Proposition 70, pp. 23–24], are the following, using the conventions introduced in Subsection A.2:
@@
It is convenient to recall that combining (17) and (19) yields [33, (85), p. 32]:
@@
In the above equations and in what follows, it is useful to recall [33, Lemma 72, p. 25]:
@@ 1.5. Higher order derivative estimates for solutions to the FRS equations
Even though our main application in the present paper is to derive complete asymptotics for the solutions constructed in [33], it is of interest to derive energy estimates using minimal assumptions. In this way, the result might be useful in other contexts. The main assumptions are the following.
@@
Remark 20.
The eigenvalues of the expansion normalised Weingarten map are not assumed to be distinct in this theorem.
@@ 1.6. Deriving asymptotics, given higher order derivative bounds
To begin with, (30a) is, e.g., an immediate consequence of [43, (269), p. 65], keeping in mind that
@@
$\mathcal{P}$ is defined by [43, (17), p. 8] and the fact that $T$ is given by (2). Next, (30b) follows from a straightforward computation and (31) follow from [39, Chapter 13].
Turning to the other direction, (1b) is an immediate consequence of (30b). Due to (31) the normal–normal and the normal–tangential components of Einstein’s equations hold. Next, combining [43, (257), p. 63] and [43, (269), p. 65] with (30a), it is clear that the spatial–spatial components of the spacetime Ricci tensor equal what they would if Einstein’s equations would hold. Combining this information with the fact that the normal–normal and the normal–tangential components of Einstein’s equations hold, it can be deduced that (1a) holds. ∎
@@
(36c) $\|N\|_{C^{k_{0}}(\Sigma)}+\|N^{-1}\|_{C^{0}(\Sigma)}+t^{1-\delta}\|\vec{e}(\ln N)\|_{C^{k_{0}}(\Sigma)}+t^{1-\delta}\|\vec{e}\varphi\|_{C^{k_{0}}(\Sigma)}$ $\leq\mathbf{c}_{0},$
(36d) $\|t\Theta-1\|_{C^{0}(\Sigma)}$ $\leq\tfrac{1}{2}$
@@ 1.7. Outline of the arguments
The main result of this article is Theorem 11. However, Theorem 11 is an immediate consequence of Theorems 19 and 24. Moreover, the latter theorems have an independent interest in their own right. In particular, it is to be expected that Theorem 24 can be used to obtain data on the singularity in other settings. For these reasons, we here focus on a description of the arguments required to prove Theorems 19 and 24.
@@
Following the results of [33], it makes sense to set as our starting point some $C^{k_{0}}$ bounds on a Fermi-Walker frame and on the components of geometric quantities in terms of this frame, along with $C^{k_{0}}$ convergence of the eigenvalues of $\mathcal{K}$, of $\Phi$ and of $\Psi$. These are precisely the estimates (36) and (35). Note that these estimates correspond to the conclusions obtained in [33] with two exceptions. First, since we do not assume the CMC condition, we need a way to control the mean curvature. This is done through the quantity $\Theta=N\theta$, which is assumed to be asymptotic to $t^{-1}$ as $t\downarrow 0$. Second, in [33], the authors obtain convergence of the lapse, $N\to 1$ in $C^{k_{0}}$ as $t\downarrow 0$. By contrast, here we only assume that $N$ is bounded in $C^{k_{0}}$, and that $N^{-1}$ is bounded in $C^{0}$. However, we point out that if the CMC condition is assumed, then the assumption that $t\Theta-1\to 0$ becomes $N-1\to 0$. We begin by deriving some basic consequences of the assumptions in Subsection 3.1.
@@
Proof of the main result. Once we have Theorems 19 and 24 at our disposal, we can combine them with the arguments in [33] in order to deduce that Theorem 11 holds. The details are to be found in Section 4.
@@ 1.8. Acknowledgements
This research was funded by the Swedish Research Council (Vetenskapsrådet), dnr. 2022-03053. The first author was supported by foundations managed by The Royal Swedish Academy of Sciences.
@@ 2.5. Estimating the lapse function
As in the proof of [33, Lemma 112, p. 52], applying $E_{\mathbf{I}}$ to (20) and then multiplying with $E_{\mathbf{I}}(N-1)$ yields
@@
(61) $\begin{split}&-\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}e_{I}(N-1)+t^{-2}\big(E_{\mathbf{I}}(N-1)\big)^{2}\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}e_{I}(N-1)[e_{I},E_{\mathbf{I}}](N-1)-[E_{\mathbf{I}},e_{I}]e_{I}(N-1)E_{\mathbf{I}}(N-1)\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}(N-1)\mathrm{div}_{h_{\mathrm{ref}}}e_{I}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ ={}&\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}\big(\gamma_{JII}e_{J}(N)-N\big\{t^{-2}-k_{IJ}k_{IJ}-e_{0}(\varphi)e_{0}(\varphi)+\tfrac{2}{n-1}V\circ\varphi\big\}\big)E_{\mathbf{I}}(N-1)\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}},\end{split}$
where we used (141). Next, note that, due to (39),